# Macaw: A Machine Code Toolbox for the Busy Binary Analyst

### v2

@@ 3.1. Reopt
Reopt (Hendrix et al., 2019) is an optimization tool that lifts functions in a compiled binary into LLVM bitcode, optimizes the intermediate bitcode, and then recompiles it into a new executable. Reopt currently supports x86-64 binaries compiled in the ELF format for Linux. The Macaw library began its life as an internal component of the Reopt tool, and much of Macaw’s design reflects considerations that were important at the time of Reopt’s development.
@@
To this end, Reopt includes a type-based function recovery pass, which is depicted in Figure 3. Starting from a list of initial entry point functions (which are either inferred from the binary or user-specified), the tool calls Macaw to discover the closure of basic blocks reachable from the currently known blocks. Once the basic blocks have been recovered (c.f. Section 2.5), Reopt performs a demand analysis, allowing it to compute the x86-64 arguments and return registers on a per-block basis, aggregating them into register “demand sets” for entire functions, which are then synthesized into appropriate argument and return values for the whole function.
@@ 3.1.1. Reopt Type Constraints
Ultimately, Reopt is concerned with distinguishing pointers from non-pointer values. Without the type inference pass, Reopt would lift all pointer arithmetic to integer arithmetic and would have to introduce an LLVM inttoptr cast every time such a value would be in address position in a memory load or store instruction. This is important because treating pointers as integers up to the last second inhibits LLVM optimizations and obfuscates pointer manipulation for both human and tool consumers of the lifted IR. Instead, pointer arithmetic can be lifted to pointer-manipulating instructions, describing them using types with richer structure, benefiting the readability of the lifted code, and enabling more complex code analyses and transformations. Knowing these types, Reopt can produce the more appropriate getelementptr LLVM instruction, operating over typed pointers. We give an instance of such an output change in Appendix B, Figure B.1.
@@ 3.2. SAW
The Software Analysis Workbench (SAW) (Dockins et al., 2016) is a verification tool that can reason about the behavior of imperative programs by proving them equivalent to functional specifications. SAW supports a variety of different source languages, including code written in C, Java, Rust, and most relevantly for this work, x86-64. SAW has been successfully used to verify a variety of industrial-strength cryptographic libraries written in a mixture of C and x86-64, including the s2n Transport Layer Security (TLS) stack (Chudnov et al., 2018), the AWS LibCrypto library (Boston et al., 2021), and the BLST library (Dodds, 2020).
@@
Here, increment_bin is the ELF binary that defines increment, and z3 is the SMT solver to discharge proof goals to during symbolic execution. This is the only command that is machine code–specific, as this is where Macaw is used to lift the assembly code in increment to a CFG and pass it on to SAW’s symbolic execution engine. Although the llvm_verify_x86 command is specific to the x86-64 ISA, it could very easily be extended to handle other architectures as well 111For example, an llvm_verify_aarch32 command is planned: anonymized for review .
@@ Acknowledgements.
- Selfridge (2019) Benjamin Selfridge. 2019. GRIFT: A richly-typed, deeply-embedded RISC-V semantics written in Haskell. In Proc. SpISA 2019: Workshop Instr. Set Architect. Specification.
@@
- Srinivasan and Reps (2015) Venkatesh Srinivasan and Thomas Reps. 2015. Synthesis of machine code from semantics. In Proceedings of the 36th ACM SIGPLAN Conference on Programming Language Design and Implementation. 596–607.
@@
Appendix A Macaw Design (Additional Details)
data MemRepr (tp :: MacawType) where
BVMemRepr
:: (1 <= w)
=> NatRepr w
-> Endianness
-> MemRepr (BVType (8*w))
FloatMemRepr
:: FloatInfoRepr f
-> Endianness
-> MemRepr (FloatType f)
PackedVecMemRepr
:: NatRepr n
-> MemRepr tp
-> MemRepr (VecType n tp)
data FloatInfoRepr (fi :: FloatInfo) where
SingleFloatRepr ::
FloatInfoRepr SingleFloat
DoubleFloatRepr ::
FloatInfoRepr DoubleFloat
data Stmt arch ids where
AssignStmt
:: Assignment arch ids tp
-> Stmt arch ids
WriteMem
:: ArchAddrValue arch ids
-> MemRepr tp
-> Value arch ids tp
-> Stmt arch ids
ExecArchStmt
:: ArchStmt arch
(Value arch ids)
-> Stmt arch ids
ArchState
:: ArchMemAddr arch)
-> RegState
(ArchReg arch)
(Value arch ids)
-> Stmt arch ids
InstructionStart
:: ArchAddrWord arch
-> Text
-> Stmt arch ids
Comment
:: Text
-> Stmt arch ids
Appendix B Reopt Case Study (Additional Details)
;; Original x86-64 instruction
cmpb $0x30,0x3(%rax) ; in AT&T syntax
cmp BYTE PTR [rax+0x3],0x30 ; in Intel syntax
;; Translation without type inference
;; %t0 corresponds to %rax and is given type i64
%t1 = add i64 %t0, 0x3
%t2 = load i8, i8* inttoptr (i64 %t1 to i8*)
%t3 = icmp eq i8 %t2, 0x30
;; Translation with type inference
;; %t0 corresponds to %rax and is given type {{i8,i8,i8},{i8,i8}}*
%t1 = getelementptr {{i8,i8,i8},{i8,i8}},
{{i8,i8,i8},{i8,i8}}* %t0, i32 0, i32 1
%t2 = getelementptr {i8,i8}, {i8,i8}* %t1, i32 0, i32 0
%t3 = load i8, i8* %t2
%t4 = icmp eq i8 %t3, 0x30

### v3

@@ 2.1.6. Symbolic Execution via Crucible
2.2. Macaw-x86: The First Macaw Dialect
Macaw-x86, which targets the x86-64 ISA, was the first ISA dialect that we developed alongside the Reopt tool (see Section 3.1). The architecture of Macaw-x86 is pictured in Figure 6. While this figure is specific to x86-64, it establishes a general pattern that we continued to use when developing future dialects, albeit with some notable differences (which we will address later).
Because Macaw-x86 was our first attempt at a dialect, we did not have an existing methodology for integrating an ISA into Macaw, nor did we have a clear picture of what design choices one could make in the process of developing a dialect. One notable decision that we made (perhaps unconsciously) during development was to hand-write most of the code in Macaw-x86. This is because the Reopt tool needed to be able to run on x86-64 binaries with a relatively quick turnaround time, and as such, we could not wait until we had 100% coverage of the x86-64 ISA before demonstrating results. As such, we implemented support for individual x86-64 instructions on an as-needed basis, manually implementing whatever scaffolding was necessary for Macaw to reason about each instruction.
2.2.1. Disassembling x86-64 Binaries Using Flexdis86
The first step that Macaw-x86 performs when analyzing a binary is to diassemble it—that is, to ingest the binary’s raw bytes into a more structured representation. A binary’s contents can be thought of as a stream of instructions, and the role of a disassembler is to parse one instruction from the stream at a time. We wrote our own x86-64 disassembler library111Flexdis86 contains both an x86-64 disassembler and assembler, but we do not use the assembler from Macaw., named Flexdis86, which represents each instruction as a Haskell data type, called an InstructionInstance. An InstructionInstance captures all of an instruction’s relevant properties, such as its name, opcode, size, and operands (be they immediate values, registers, addresses, or otherwise).
The main entrypoint into Flexdis86 is the disassembleInstruction function:
The ByteReader class characterizes Haskell monads that can represent instruction streams. Flexdis86 is primarily meant to be used with a ByteReader instance that efficiently reads a binary from disk, but Flexdis86 also supports a simpler ByteReader instance that can be used to mock the contents of a binary for testing purposes.
Because x86-64 is a very large ISA with well over 1,000 instructions, writing a feature-complete disassembler for x86-64 is no small feat. To make this task more manageable, disassembleFunction takes an NextOpcodeTable argument, which represents a lookup table that maps an instruction opcode to possible InstructionInstance values:
To understand how NextOpcodeTable works, consider the pop instruction. Some versions of the pop instruction have opcodes that begin with the byte 0f, so a NextOpcodeTable would have an entry at index 0f containing a OpcodeTable that handles the remaining opcode bytes. This OpcodeTable would then contain another, smaller NextOpcodeTable with entries at indices a1 and a9, each containing OpcodeTableEntry values for pop. These table entries represent two variants of the instruction with the opcodes 0f a1 and 0f a9, respectively. The smaller NextOpcodeTable would also map any opcode bytes for other instructions whose opcodes begin with 0f (e.g., the addps instruction, which can use the opcode 0f 58).
Constructing a NextOpcodeTable is by far the most tedious part of Flexdis86’s disassembler. To make this process less tedious, we partially automate it by leveraging the udis86 project (Thampi, 2014), which implements an x86 and x86-64 disassembler for C. Flexdis86 uses udis86’s XML table describing all x86-64 instructions to automatically generate a NextOpcodeTable value. This automation does save us quite a bit of time, but it is not perfect. Flexdis86’s version of the XML table makes some manual tweaks, such as adding additional instructions not covered by udis86. We also do not automate any of the logic for turning InstructionCandidates into full-fledged InstructionInstances, which can be surprisingly complicated due to the sheer number of instruction prefixes that x86-64 uses to distinguish different instructions.
2.2.2. Encoding x86-64 Semantics Using Macaw-x86
After disassembling a binary’s instructions, the next step is to reify them into a semantic representation that Macaw can reason about. The Macaw-x86 dialect accomplishes this by defining an InstructionDef for each x86-64 instruction, which converts an InstructionInstance from Flexdis86 into a list of Macaw basic blocks:
We also define a small domain-specific language for defining InstructionDef values that makes it more obvious what an instruction’s semantics mean at a glance. As a small example, here is the complete semantics for the cbw (convert byte to word) instruction in Macaw-x86:
These semantics use defNullary to signify that the cbw instruction does not take any operands. It first reads a byte from the al register (get al), uses sign extension (sext) to convert the byte into a two-byte value (n16), and finally copies the two-byte value into the ax register (ax .= ...). The InstructionDef DSL also includes shorthand for other common machine code idioms (immediate values, memory locations, offsets, etc.) that are not shown above.
Unlike Flexdis86, where we made an effort to automate at least some parts of the implementation, all of the semantics implemented in Macaw-x86 are completely hand-written. This was largely a practical consideration, as we wanted to demonstrate results for the subset of x86-64 instructions that we encountered the most on the Reopt project, and hand-writing the semantics for these hand-picked instructions proved a fast way to accomplish this. Our methodology was to carefully read the Intel x86-64 ISA manual (Intel, 2024) and extract formal semantics for each instruction’s entry in the manual. Although the list of instructions covered by Macaw-x86’s semantics has grown over the years, it is still far from complete.
The choice to hand-write Macaw-x86’s semantics, while expedient in the short-term, poses a challenge to long-term maintenance. If we use Macaw-x86 on a binary containing an instruction that is not covered by the list of covered instructions, then we must add the instruction’s semantics manually, which is a time-consuming process. Not only that, but this is an error-prone process, as human programmers are liable to make mistakes when translating the ISA manual’s prose.
2.3. Macaw-PPC: A More Automated Dialect
As we gained experience developing the Macaw-x86 dialect, it became increasingly clear that manually curating dialects for large ISAs did not scale well as well as we’d like. We decided that if we were to add another Macaw dialect for a big ISA, we would make a concerted effort to automate as much of the process as was reasonable. That opportunity came when we developed the next dialect, Macaw-PPC. The architecture of Macaw-PPC is pictured in Figure 7. Our focus with Macaw-PPC was to support both 32- and 64-bit PowerPC binaries, although we tend to exercise the 32-bit PowerPC code paths more often due to its prevalence in embedded software.
2.3.1. Disassembling PowerPC Binaries Using Dismantle
The first step in developing Macaw-PPC was to generate a disassembler. We used the udis86 disassembler in order to (partially) generate Macaw-x86’s disassembler, but because udis86 is specific is x86, we could not do something similar PowerPC. As a result, we developed a more general solution: a cross-architecture assembler and diassembler generation library called Dismantle. Dismantle would become the basis for Macaw-PPC’s disassembler, as well as the the disassembler for most Macaw dialects going forward.
Dismantle works by consulting the data provided by LLVM’s TableGen tool (LLVM Project, 2024). LLVM uses TableGen as part of its code generator to describe the encodings, types, and operands of machine instructions. TableGen data is also ideal for generating disassemblers, so Dismantle is able to use TableGen with little or no changes required.222Typically, the only time where we need to modify the TableGen data for Dismantle’s purposes is to work around LLVM bugs.
TableGen data files are large, and we don’t want to have to load a data file every time we disassemble a PowerPC binary, as this would be prohibitively expensive. Our solution is to instead load the data file only once when compiling the package. We use Template Haskell (Sheard and Jones, 2002) to load the PowerPC data file at compile time, parse it, and then use Template Haskell to generate code for an efficient disassembler based on the TableGen information. Note that we trade runtime efficiency in exchange for much more work at compile-time: it can take several minutes to compile all of Dismantle. This is a theme that would continue to reappear as we moved in the direction of increased automation.
2.3.2. Semi-Automated PowerPC Semantics Using SemMC
Just like with disassembling PowerPC, we sought to futher automate the process of encoding the semantics of PowerPC instructions into Macaw. Luckily for us, Heule et al. (2016) had recently published work on how to automatically generate formal semantics for x86-64using a techique called stratified synthesis. Although this work came out well after we hand-wrote our own formal semantics for x86-64, we decided to apply stratified synthesis when developing PowerPC semantics. We developed a library for performing this form of synthesis (for PowerPC and for other ISAs) and named it SemMC (Semantics of Machine Code).
The SemMC library is designed to learn machine code semantics by starting with a core base set of instructions for which the semantics are hand-written. Using this base set, we execute SemMC on a series of fuzzer-generated machine states and instructions and use the results to learn the semantics of instructions that are not in the base set. To further increase our confidence that the learned semantics are correct, we run the randomly generated machine states on actual PowerPC hardware and use that as an oracle for what behavior should be considered correct. We repeat this process until we have a complete formal description of the ISA’s semantics. In this sense, SemMC’s stratified synthesis is semi-automated: some manual effort is required to describe the semantics of the base set, but everything else is fully automated.
Much like with the semantics implementation in Macaw-x86, the semantics of SemMC base set instructions are specified using a DSL. For example, here are the semantics of PowerPC’s nand instruction in SemMC:
Here, definePPCOpcode takes the instruction name and what instruction form to use.333Instruction forms describe how instructions are encoded in PowerPC. It is not essential to understand how instruction forms work to read this section. We use xform3c, which indicates an X-form instruction with three operands. These operands are then bound as rA, rB, and rC, which can then be used in the instruction’s definition. Finally, we use defLoc to store the result of performing a bitwise-AND of the values from the rS and rB operands into the location that the operand rA uses.
SemMC converts the DSL into an intermediate language called What4. What4 programs are stored on disk using an S-expression–based representation so that upstream programs (e.g., Macaw-PPC) can parse and interpret the semantics they contain. For example, the What4 representation of the nand semantics above is the following S-expression:
The What4 representation of SemMC semantics explicitly indicate which operands are inputs or outputs (via the in section), and it precisely captures how each output is modified when invoking the instruction. In addition to updating the rA operand, the What4 version also updates the instruction pointer IP, which is a lower-level detail that was elided in the DSL encoding.
The DSL is a convenient way for humans to write What4 semantics in the base set, but SemMC also needs to be able to synthesize What4 semantics based on fuzzing. This is a more complicated process, as SemMC must be able to reason about how a machine’s state updates when invoking an instruction, and it must be able to do this where not all parts of the machine state are known ahead of time—that is, when some parts of the machine state are symbolic.
To make this process simpler, SemMC leverages the Crucible (Christiansen et al., 2019) symbolic execution library. Crucible is primarily designed for imperative programming languages, such as C or Rust, but Crucible also works well for machine code as well. SemMC leverages Crucible when synthesizing What4 semantics for instructions not in the base set.
Finally, once we have What4 semantics for all PowerPC instructions, we translate the What4 programs into Macaw CFGs using the Macaw-SemMC library. As was the case with Dismantle, we do not want to perform all of this conversion at runtime, as we would need to read many, many What4 S-expression files in order to look up the semantics for all of the instructions used in a typical PowerPC program. As such, Macaw-SemMC uses Template Haskell to look up all What4 S-expression files at compile time and generate Macaw CFG–based semantics for each PowerPC instruction, thereby avoiding the need to consult any What4 files at runtime.
The payoff of introducing SemMC as part of the semantics translation is that we gain a significant amount of automation in the overall process, thereby avoiding some of the maintenance headaches that pervade Macaw-x86’s fully manual approach to semantics. There is a cost to doing so, however: SemMC’s extra layers of abstraction make it so that it approximately two minutes to compile the semantics for the entire PowerPC ISA.
2.4. Macaw-AArch32: An Even More Automated Dialect
Using SemMC to automate the development of Macaw-PPC offered significant advantages in how confident we were in the resulting PowerPC semantics’ correctness, and it made the code more maintainable as well. The SemMC approach is not foolproof, however. SemMC still requires human beings to curate semantics for base set instructions by hand, and while the size of the base set is smaller than the size of the overall ISA, curating the base set can still be a tedious (and potentially error-prone) process.
When developing Macaw-AArch32, a Macaw dialect for the 32-bit ARM ISA, we had an opportunity to develop the level of automation even further. ARM maintains machine-readable, executable specifications of every instruction (Reid, 2016), which gives Macaw-AArch32 an official source of truth for how to disassemble and interpret any given instruction. Combining the offical ARM specifications with Template Haskell allows us to automate almost every aspect of Macaw-AArch32’s development.
The architecture of Macaw-AArch32 is pictured in Figure 8. Like Macaw-PPC, Macaw-AArch32also relies on the Dismantle package in order to disassemble binaries. One notable difference in how Dismantle’s PowerPC and AArch32 dialects work is that Dismantle uses LLVM TableGen data for PowerPC, whereas it uses the ARM XML specifications for AArch32. Note that the constructing parse tables via the XML specifications is quite slow: it takes over an hour to build tables for the entire ISA. As such, we pre-compute the parse tables and store them on disk so that they can simply be loaded directly during subsequent rebuilds of Dismantle.
There is also a SemMC dialect for AArch32, but unlike the SemMC dialect for PowerPC, it does not generate semantics for ARM instructions using stratified synthesis. Instead, it uses a dedicated library (named ASL-Translator) to convert the ARM XML specification’s semantics directly into the What4 intermediate language that SemMC uses. The XML specification is written in ASL (Reid, 2016), a custom-purpose imperative programming language with a strong types. As it turns out, the typing discipline that ASL uses is a natural fit for Crucible, so ASL-Translatorheavily relies on Crucible to translate the imperative, ASL-based semantics into a purely functional, What4-based encoding.
Note that ASL-Translator is almost, but not completely, automated. ASL-Translator manually defines a handful of primitives to give semantics to low-level ASL operations (e.g., bitvector slicing, querying the program counter value, etc.) This is by far the trickiest part of ASL-Translator to write, and the vast majority of bug reports ultimately step from incorrectly implementing these primitives.
Like Macaw-PPC, Macaw-AArch32’s heavy use of Template Haskell (both in ASL-Translatorand in other parts of the dialect) incurs a significant compile-time performance penalty. In fact, the performance penalty is even more severe for Macaw-AArch32: it takes well over 5 minutes to compile the parts of the dialect that rely on ASL-Translator-generated semantics. While it is likely that these times could be improved by working harder to reduce the size of the generated code, this reflects the general trend that the more one uses Template Haskell, the longer one has to wait.
2.5. Macaw-RISCV: Revisiting a Manually-Written Dialect
The most recently introduced Macaw dialect is Macaw-RISCV, which targets the RISC-V architecture. The architecture of Macaw-RISCV is pictured in Figure 9. In stark contrast to previous Macaw dialects, which have moved increasingly in the direction of automatically generating disassemblers and semantics for their respective ISAs, Macaw-RISCV implements its RISC-V disassembler and semantics completely by hand. This choice was motivated by the fact that RISC-V adheres to the reduced instruction set computer (RISC) philosophy. The base set of RISC-V instructions (not including extensions) only consists of 48 instructions, which are composed of a relatively small number of different opcodes. As such, it is viable to hand-write a Macaw dialect for RISC-V.
2.6. The Manual-Automated Spectrum
Each Macaw dialect exists on a spectrum of how much automation is used to implement the code supporting the dialect’s ISA, as depicted in Figure 10. One far end of the spectrum is hand-writing all of the code (e.g., Macaw-RISCV). The opposite end of the spectrum is fully automating the generation of all ISA-related code (no dialect meets this criterion exactly, but Macaw-AArch32 comes the closest).
Having developed four Macaw dialects, we have honed some intuition regarding whether to use a hand-written approach, an automated approach, or somewhere in between when creating a new dialect. There are various questions one must ask to determine which approach is right for a given ISA:
How many different instructions are there? It is possible to exhaustively go through all of the instructions in a small, RISC architecture, but for non-RISC architectures, the pain of going through every instruction will be more pronounced.
How clear is the ISA specification regarding instruction semantics? Different ISA semantics are written with different levels of formality. For instance, the ARM XML specification is written with extreme formality, while the Intel x86-64 ISA manual is less so, sometimes requiring humans to interpret dense English-language prose in order to determine how certain instructions behave. For the latter case, a SemMC-like approach can help.
How tolerant are you of long compile times? While automatically generating code with Template Haskell can increase ISA coverage and confidence in the generated code, it comes with a severe downside of noticeably increasing compile times. It is worth asking the question of whether this compile-time slowdown is acceptable for a given project.
@@ 3.1. Reopt
Reopt (Hendrix et al., 2019) is an optimization tool that lifts functions in a compiled binary into LLVM bitcode, optimizes the intermediate bitcode, and then recompiles it into a new executable. Reopt currently supports x86-64 binaries compiled in the ELF format for Linux. The Macaw library began its life as an internal component of the Reopt tool, and much of Macaw’s design reflects considerations that were important at the time of Reopt’s development. The source code for Reopt is publicly available at https://github.com/GaloisInc/reopt.
@@
To this end, Reopt includes a type-based function recovery pass, which is depicted in Figure 11. Starting from a list of initial entry point functions (which are either inferred from the binary or user-specified), the tool calls Macaw to discover the closure of basic blocks reachable from the currently known blocks. Once the basic blocks have been recovered (c.f. Section 2.1.5), Reopt performs a demand analysis, allowing it to compute the x86-64 arguments and return registers on a per-block basis, aggregating them into register “demand sets” for entire functions, which are then synthesized into appropriate argument and return values for the whole function.
@@ 3.1.1. Reopt Type Constraints
Ultimately, Reopt is concerned with distinguishing pointers from non-pointer values. Without the type inference pass, Reopt would lift all pointer arithmetic to integer arithmetic and would have to introduce an LLVM inttoptr cast every time such a value would be in address position in a memory load or store instruction. This is important because treating pointers as integers up to the last second inhibits LLVM optimizations and obfuscates pointer manipulation for both human and tool consumers of the lifted IR. Instead, pointer arithmetic can be lifted to pointer-manipulating instructions, describing them using types with richer structure, benefiting the readability of the lifted code, and enabling more complex code analyses and transformations. Knowing these types, Reopt can produce the more appropriate getelementptr LLVM instruction, operating over typed pointers. We give an instance of such an output change in Figure 12.
;; Original x86-64 instruction
cmpb $0x30,0x3(%rax) ; in AT&T syntax
cmp BYTE PTR [rax+0x3],0x30 ; in Intel syntax
;; Translation without type inference
;; %t0 corresponds to %rax and is given type i64
%t1 = add i64 %t0, 0x3
%t2 = load i8, i8* inttoptr (i64 %t1 to i8*)
%t3 = icmp eq i8 %t2, 0x30
;; Translation with type inference
;; %t0 corresponds to %rax and is given type {{i8,i8,i8},{i8,i8}}*
%t1 = getelementptr {{i8,i8,i8},{i8,i8}},
{{i8,i8,i8},{i8,i8}}* %t0, i32 0, i32 1
%t2 = getelementptr {i8,i8}, {i8,i8}* %t1, i32 0, i32 0
%t3 = load i8, i8* %t2
%t4 = icmp eq i8 %t3, 0x30
@@ 3.2. SAW
The Software Analysis Workbench (SAW) (Dockins et al., 2016) is a verification tool that can reason about the behavior of imperative programs by proving them equivalent to functional specifications. SAW supports a variety of different source languages, including code written in C, Java, Rust, and most relevantly for this work, x86-64. SAW has been successfully used to verify a variety of industrial-strength cryptographic libraries written in a mixture of C and x86-64, including the s2n Transport Layer Security (TLS) stack (Chudnov et al., 2018), the AWS LibCrypto library (Boston et al., 2021), and the BLST library (Dodds, 2020). The source code for SAW is publicly available at https://github.com/GaloisInc/saw-script.
@@
Here, increment_bin is the ELF binary that defines increment, and z3 is the SMT solver to discharge proof goals to during symbolic execution. This is the only command that is machine code–specific, as this is where Macaw is used to lift the assembly code in increment to a CFG and pass it on to SAW’s symbolic execution engine. Although the llvm_verify_x86 command is specific to the x86-64 ISA, it could very easily be extended to handle other architectures as well 444For example, an llvm_verify_aarch32 command is planned: anonymized for review .
@@ 3.3. Other Case Studies
The source code for Ambient-Verifier is publicly available at https://github.com/GaloisInc/ambient-verifier.
@@
The source code for PATE is publicly available at https://github.com/GaloisInc/pate.
@@
The source code for Cerridwen is publicly available at https://github.com/GaloisInc/cerridwen.
@@
The source code for Renovate is publicly available at https://github.com/GaloisInc/renovate.
@@
The source code for MCTrace is publicly available at https://github.com/GaloisInc/mctrace.
@@
The source code for Surveyor is publicly available at https://github.com/GaloisInc/surveyor.
@@ Acknowledgements.
- Heule et al. (2016) Stefan Heule, Eric Schkufza, Rahul Sharma, and Alex Aiken. 2016. Stratified synthesis: automatically learning the x86-64 instruction set. In Proceedings of the 37th ACM SIGPLAN Conference on Programming Language Design and Implementation (Santa Barbara, CA, USA) (PLDI ’16). Association for Computing Machinery, New York, NY, USA, 237–250. https://doi.org/10.1145/2908080.2908121
- Intel (2024) Intel. 2024. Intel 64 and IA-32 Architectures Software Developer’s Manual. https://www.intel.com/content/www/us/en/developer/articles/technical/intel-sdm.html
@@
- LLVM Project (2024) LLVM Project. 2024. TableGen Overview. https://llvm.org/docs/TableGen/
@@
- Sheard and Jones (2002) Tim Sheard and Simon Peyton Jones. 2002. Template meta-programming for Haskell. In Proceedings of the 2002 ACM SIGPLAN Workshop on Haskell (Pittsburgh, Pennsylvania) (Haskell ’02). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/581690.581691
@@
- Thampi (2014) Vivek Thampi. 2014. udis86: Disassembler Library for x86 and x86-64. https://github.com/vmt/udis86